Session: May 1999 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Sat May 29 08:05:43 METDST 1999) Success
No known problems

Onsala (A. Polatidis at Sat May 29 16:55:50 METDST 1999) Success
WE have problems with the head to tape contact, particularly at double speed. This resluts in a reduction in the recording quality, causing very high parity errors. Most of the data should be recoverable. We are working to solve this problem with consultation from Haystack.

Westerbork (Tony Foley at Mon May 31 15:41:25 METDST 1999) Success
No known problems.Used 12 telescopes in tied-array

Noto (C.Trigilio at Mon May 31 16:47:45 METDST 1999) Success

Effelsberg (A. Kraus at Tue Jun 1 08:48:06 METDST 1999) Success

Jodrell (Alastair Gunn at Tue Jun 1 15:53:00 METDST 1999) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Fri Jun 18 19:32:10 METDST 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may99 directory under the name: w018f.antab. Read the comments in the file to get further information.